### v2

# The Revolution of Multimodal Large Language Models: A Survey

Abstract
Connecting text and visual modalities plays an essential role in generative intelligence. For this reason, inspired by the success of large language models, significant research efforts are being devoted to the development of Multimodal Large Language Models (MLLMs). These models can seamlessly integrate visual and textual modalities, while providing a dialogue-based interface and instruction-following capabilities. In this paper, we provide a comprehensive review of recent visual-based MLLMs, analyzing their architectural choices, multimodal alignment strategies, and training techniques. We also conduct a detailed analysis of these models across a wide range of tasks, including visual grounding, image generation and editing, visual understanding, and domain-specific applications. Additionally, we compile and describe training datasets and evaluation benchmarks, conducting comparisons among existing models in terms of performance and computational requirements. Overall, this survey offers a comprehensive overview of the current state of the art, laying the groundwork for future MLLMs.
The Revolution of Multimodal Large Language Models: A Survey
Davide Caffagni1111Equal contribution., Federico Cocchi1,2111Equal contribution., Luca Barsellotti1111Equal contribution., Nicholas Moratelli1111Equal contribution., Sara Sarto1111Equal contribution., Lorenzo Baraldi2111Equal contribution., Lorenzo Baraldi1, Marcella Cornia1,  and Rita Cucchiara1,3 1University of Modena and Reggio Emilia, Italy 2University of Pisa, Italy 3IIT-CNR, Italy 1{name.surname}@unimore.it2{name.surname}@phd.unipi.it
1 Introduction
The introduction of the attention operation and the Transformer architecture Vaswani et al. (2017) has enabled the creation of models capable of handling various modalities on an increasingly large scale. This advancement is largely attributed to the versatility of the operator and the adaptability of the architecture. Initially, this breakthrough was leveraged for language-specific models Devlin et al. (2018); Brown et al. (2020) but quickly extended to support diverse modalities Li et al. (2019); Lu et al. (2019) and facilitate their integration within unified embedding spaces Radford et al. (2021).
Figure 1: General architecture of Multimodal Large Language Models (MLLMs), composed of a visual encoder, a language model, and an adapter module that connects visual inputs to the textual space.
The surge in sophisticated Large Language Models (LLMs), particularly their capacity for in-context learning, has encouraged researchers to broaden the scope of these models to encompass multiple modalities, both as inputs and outputs. This expansion has led to the development of cutting-edge models such as GPT-4V Achiam et al. (2023) and Gemini Anil et al. (2023), showcasing state-of-the-art performance.
The development of Multimodal Large Language Models (MLLMs) entails merging single-modality architectures for vision and language, establishing effective connections between them through vision-to-language adapters, and devising innovative training approaches. These methodologies are crucial for ensuring modality alignment and the ability to follow instructions accurately.
In a context marked by the rapid release of new models, our goal is to offer an exhaustive overview of the MLLM landscape, with a focus on models exploiting the visual modality. This overview serves as both an update on the current state and a source of inspiration for future developments. We identify three core aspects that define these models: their architecture, training methodologies, and the tasks they are designed to perform. We begin by detailing the prevalent choices for vision encoders and adapter modules that equip LLMs with cross-modal capabilities. Following this, we delve into the training processes and data utilized. We then explore the range of tasks addressed by MLLMs. The review concludes with a discussion of the persisting challenges in the field and the promising directions for future research. Further details on training data, evaluation datasets, performance and computational requirements are reported in the supplementary material.
The motivation behind this survey stems from an emerging scientific interest in the field of MLLMs, as evidenced by the constant increase in published works. In comparison with existing surveys on the topic Yin et al. (2023a); Wu et al. (2023b); Huang et al. (2023a), our paper exhibits substantial differences. Notably, it addresses several critical areas that were overlooked in prior works, including visual grounding, image generation, and editing. Furthermore, our survey details the main components utilized by each discussed MLLM, such as the visual encoders and the specific LLM employed. Additionally, our analysis offers a comparative perspective on the performance and hardware requirements of the discussed papers, incorporating both quantitative results and detailed information on benchmarks. Through this comprehensive approach, our survey aims to fill the existing gaps and provide a more nuanced understanding of the current landscape in the field.
2 Empowering LLMs with Multimodal Capabilities
2.1 Preliminaries
Large Language Models. Brown et al. (2020) discovered that in-context learning, i.e., prepending the prompt with a few examples demonstrating the desired output of an LLM Chowdhery et al. (2023); Hoffmann et al. (2022); Tay et al. (2022), improves its performance, especially over unseen tasks. Generalization can be further enhanced by providing the LLM with the natural language description of the desired task for each training sample. This technique, called instruction-tuning Chung et al. (2022); Wang et al. (2022b, a); Jiang et al. (2024), turns out to be critical for aligning the behavior of an LLM with that of humans and currently empowers the most advanced LLMs, eventually boosted via reinforcement learning from human feedback (RLHF) Ouyang et al. (2022); Achiam et al. (2023); Chen et al. (2023l); Bai et al. (2023a).
PEFT. When a pre-trained LLM needs to be adapted to a specific domain or application, parameter-efficient fine-tuning (PEFT) schemes represent an important alternative to train the entire LLM, since these strategies only introduce a few new parameters. Among these, prompt-tuning Hambardzumyan et al. (2021); Lester et al. (2021); Li and Liang (2021); Liu et al. (2023j) learns a small set of vectors to be fed to the model as soft prompts before the input text. Differently, LoRA Hu et al. (2021) constrains the number of new weights by learning low-rank matrices. This technique is orthogonal to quantization methods such as QLoRA (Dettmers et al., 2024), which further decreases the memory footprint of the LLM compared to the usual half-precision weights.
Towards Multimodal LLMs. The development of MLLMs follows a similar path to that of LLMs, with Flamingo Alayrac et al. (2022) being the first to explore in-context learning at scale in the vision-language field. Then, visual instruction-tuning Liu et al. (2023e) quickly became the most prominent training paradigm also in the multimodal domain, as well as the use of PEFT techniques to fine-tune the LLM. Any MLLM contains at least three components (Fig. 1): an LLM backbone serving as an interface with the user, one (or more) visual encoders, and one or more vision-to-language adapter modules. Popular choices for the LLM backbone often fall into the LLaMA family Touvron et al. (2023a, b), given that their weights are freely accessible, they have been trained on public data solely, and they boast different sizes to accommodate various use cases. In addition, their derivative versions are popular as well, such as Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023). The former fine-tunes LLaMA on instructions written using GPT-3, while the latter exploits user-shared conversations with ChatGPT OpenAI (2022). Alternatives are OPT Zhang et al. (2022b), Magneto Wang et al. (2023b), MPT MosaicML (2023), and the instruction-tuned Chung et al. (2022) or multilingual Xue et al. (2020) flavors of T5 Raffel et al. (2020), an encoder-decoder language model pre-trained for multiple tasks.
Pre-Training of Model Components. The main components of MLLMs are the visual encoder and the language model. The visual encoder is designed to provide LLMs with visual information and the most used ones are CLIP-based architectures Radford et al. (2021); Wortsman et al. (2022) whose pre-training objective is the alignment between CLIP embeddings, obtained thanks to a contrastive loss that aligns the correct image-text pairs. An exception is the EVA-CLIP models family Fang et al. (2023), which exploits a MAE pre-training strategy He et al. (2022) to reconstruct the masked-out image-text aligned visual features, conditioned on visible image patches. On the other hand, LLMs primarily rely on the widely employed Transformer model, although the Mamba architecture Gu and Dao (2023) has also emerged in recent times. This proposes to make a State-Space Model (SSM) time-dependent, effectively creating a selective SSM with favorable properties: (i) inference costs and memory requirements that scale linearly with the sequence length, and (ii) efficient parallel training thanks to a smart GPU implementation of the algorithm. Similar to Transformers, Mamba models for language modeling are pre-trained using the next token prediction task. Very recent studies propose MLLMs featuring Mamba as the language backbone Qiao et al. (2024); Zhao et al. (2024).
A summary of the MLLMs covered in this survey is reported in Table 1, indicating for each model the LLM on which it is based, the visual encoder, the adapter used to connect visual and language components, whether the MLLM is trained with visual instruction tuning or not, and a short list of the main tasks and capabilities.
Visual V2L VInstr.
Model LLM Encoder Adapter Tuning Main Tasks & Capabilities
BLIP-2 Li et al. (2023g) FlanT5-XXL-11B EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning, Retrieval
FROMAGe Koh et al. (2023b) OPT-6.7B CLIP ViT-L Linear Visual Dialogue, Captioning, Retrieval
Kosmos-1 Huang et al. (2023b) Magneto-1.3B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
LLaMA-Adapter V2 Gao et al. (2023) LLaMA-7B CLIP ViT-L Linear VQA, Captioning
OpenFlamingo Awadalla et al. (2023) MPT-7B CLIP ViT-L XAttn LLM VQA, Captioning
Flamingo Alayrac et al. (2022) Chinchilla-70B NFNet-F6 XAttn LLM Visual Dialogue, VQA, Captioning
PaLI Chen et al. (2023j) mT5-XXL-13B ViT-e XAttn LLM Multilingual, VQA, Captioning, Retrieval
PaLI-X Chen et al. (2023h) UL2-32B ViT-22B XAttn LLM Multilingual, VQA, Captioning
LLaVA Liu et al. (2023e) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
MiniGPT-4 Zhu et al. (2023a) Vicuna-13B EVA ViT-g Linear VQA, Captioning
mPLUG-Owl Ye et al. (2023c) LLaMA-7B CLIP ViT-L Q-Former Visual Dialogue, VQA
InstructBLIP Dai et al. (2023) Vicuna-13B EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning
MultiModal-GPT Gong et al. (2023) LLaMA-7B CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
LaVIN Luo et al. (2023) LLaMA-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
Otter Li et al. (2023b) LLaMA-7B CLIP ViT-L XAttn LLM VQA, Captioning
Kosmos-2 Peng et al. (2023) Magneto-1.3B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning, Referring, REC
Shikra Chen et al. (2023f) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
Clever Flamingo Chen et al. (2023b) LLaMA-7B CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
SVIT Zhao et al. (2023a) Vicuna-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
BLIVA Hu et al. (2024) Vicuna-7B EVA ViT-g Q-Former+Linear Visual Dialogue, VQA, Captioning
IDEFICS Laurençon et al. (2024) LLaMA-65B OpenCLIP ViT-H XAttn LLM Visual Dialogue, VQA, Captioning
Qwen-VL Bai et al. (2023b) Qwen-7B OpenCLIP ViT-bigG Q-Former Visual Dialogue, Multilingual, VQA, Captioning, REC
StableLLaVA Li et al. (2023i) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
Ferret You et al. (2023) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, Captioning, Referring, REC, GroundCap
LLaVA-1.5 Liu et al. (2023d) Vicuna-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
MiniGPT-v2 Chen et al. (2023e) LLaMA-2-7B EVA ViT-g Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
Pink Xuan et al. (2023) Vicuna-7B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
CogVLM Wang et al. (2023c) Vicuna-7B EVA ViT-E MLP Visual Dialogue, VQA, Captioning, REC
DRESS Chen et al. (2023l) Vicuna-13B EVA ViT-g Linear Visual Dialogue, VQA, Captioning
LION Chen et al. (2023d) FlanT5-XXL-11B EVA ViT-g Q-Former+MLP Visual Dialogue, VQA, Captioning, REC
mPLUG-Owl2 Ye et al. (2023d) LLaMA-2-7B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
SPHINX Lin et al. (2023b) LLaMA-2-13B Mixture Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
Honeybee Cha et al. (2023) Vicuna-13B CLIP ViT-L ResNet blocks Visual Dialogue, VQA, Captioning
VILA Lin et al. (2023a) LLaMA-2-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
SPHINX-X Gao et al. (2024) Mixtral-8$\times$7B Mixture Linear Visual Dialogue, Multilingual, VQA, Captioning, Referring, REC
Table 1: Summary of generalist MLLMs for vision-to-language tasks. For each model, we indicate the LLM used in its best configuration as shown in the original paper ($\lozenge$: LLM training from scratch; $\blacklozenge$: LLM fine-tuning; $\blacktriangle$: LLM fine-tuning with PEFT techniques; $\bigstar$: frozen LLM). The ✻ marker indicates variants to the reported vision-to-language adapter, while gray color indicates models not publicly available.
2.2 Visual Encoder
In MLLMs, one of the key components is a visual encoder, which is specifically designed to provide the LLM with the visual extracted features. It is common to employ a frozen pre-trained visual encoder while training only a learnable interface that connects visual features with the underlying LLM. While this is usually done using low-resolution images with fixed aspect ratios, some attempts Xu et al. (2024); Li et al. (2023l) involve adapting pre-trained visual backbones to handle images of different resolutions and aspect ratios. Further details on how to handle higher-resolution images are provided in the supplementary.
The most often employed visual encoders are based on pre-trained Vision Transformer (ViT) models with a CLIP-based objective to exploit the inherent alignment of CLIP embeddings. Popular choices are the ViT-L model from CLIP Radford et al. (2021), the ViT-H backbone from OpenCLIP Wortsman et al. (2022), and the ViT-g version from EVA-CLIP Fang et al. (2023).
As shown in Li et al. (2023g), a stronger image encoder leads to better performance. Building on this insight, Lin et al. (2023b) and Gao et al. (2024) propose an ensemble of frozen visual backbones to capture robust visual representations and different levels of information granularity. Concurrently, PaLI models Chen et al. (2023j, h), noticing an imbalance between language and visual parameters, propose scaling the visual backbone respectively to a 4- and 22-billion parameter ViT.
The utilization of such large and powerful models is made feasible by the common practice of maintaining the visual encoder frozen during training, as observed in Li et al. (2023g); Huang et al. (2023b); Gao et al. (2023); Chen et al. (2023f). However, employing a frozen visual encoder has some limitations, primarily due to the constrained number of parameters, resulting in an inadequate alignment between the visual and language modalities. Specifically, the dense features, extracted from the visual model, may fragment the fine-grained image information and bring large computation due to the lengthy sequence when fed into the language model. To mitigate this issue, other approaches Ye et al. (2023c, d) employ a two-stage training paradigm. In the first stage, they incorporate a trainable visual backbone while maintaining the pre-trained LLM frozen. According to their findings, enabling the vision encoder to be trainable enhances performance on tasks such as visual question answering or visual description. However, it may lead to performance degradation in other tasks, indicating a degree of forgetting and damage to the general visual representation.
2.3 Vision-to-Language Adapters
The simultaneous presence of inputs from different modalities emphasizes the need to incorporate a module capable of delineating latent correspondences within these unimodal domains. These modules, termed as “adapters”, are intended to facilitate interoperability between the visual and textual domains. A spectrum of different adapters are used in common MLLMs, ranging from elementary architectures such as linear layers or MLP to advanced methodologies such as Transformer-based solutions, exemplified by the Q-Former model, and conditioned cross-attention layers added to the LLM.
Linear and MLP Projections. The most straightforward approach for projecting visual inputs into textual embeddings involves learning a linear mapping, which translates visual features to the same dimensionality as the textual counterpart. Some approaches like LLaMA-Adapter Gao et al. (2023) and FROMAGe Koh et al. (2023b) only employ a single linear layer to perform the multimodal connection, while LLaVA-1.5 Liu et al. (2023d) adopts a two-layer MLP, showing improved multimodal capabilities. Despite its widespread adoption in early MLLMs, the use of linear projections has proven highly effective even in recent methods with a more advanced understanding of the visual input Chen et al. (2023f); Lin et al. (2023a); Wang et al. (2023c); You et al. (2023); Zhao et al. (2023a). It is, therefore, a simple yet effective technique for aligning visual features with textual counterparts. A different approach Cha et al. (2023) proposes to replace linear layers with convolutional ones, demonstrating moderate improvements.
Q-Former. It is a Transformer-based model proposed in BLIP-2 Li et al. (2023g) and then used in several other approaches Chen et al. (2023d); Dai et al. (2023); Hu et al. (2024). It is characterized by its adaptable architecture, which consists of two Transformer blocks that share mutual self-attention layers, facilitating the alignment process between visual and textual representations. It involves a set of learnable queries that interact within the self-attention layers and interface with visual features via a cross-attention mechanism. Textual and visual elements communicate via shared self-attention within the modules.
Drawing inspiration from the Q-Former, various modified versions have been introduced. In this regard, mPLUG-Owl models Ye et al. (2023c, d) simplify the Q-Former architecture and propose a visual abstractor component that operates by condensing visual information into distinct learnable tokens to derive more semantically enriched visual representations. On the same line, Qwen-VL Bai et al. (2023b) compresses visual features using a single-layer cross-attention module with learnable queries also incorporating 2D positional encodings.
Additional Cross-Attention Layers. This approach has been proposed in Flamingo Alayrac et al. (2022) with the integration of dense cross-attention blocks among the existing pre-trained layers of the LLM. The newly added layers are often combined with a zero-initialized tanh-gating mechanism to ensure that, upon initialization, the conditioned model acts as its original version. The use of additional cross-attention layers imposes the need to train them from scratch, increasing the number of trainable parameters compared to other alternatives. To reduce computational complexity, this strategy is usually paired with a Perceiver-based component Jaegle et al. (2021) that reduces the number of visual tokens before they are fed to the LLM. Since its introduction, several models Awadalla et al. (2023); Chen et al. (2023b); Laurençon et al. (2024); Li et al. (2023b) employ this technique to connect the visual modality with the underlying LLM, demonstrating enhanced training stability and improved performance.
2.4 Multimodal Training
Starting from a pre-trained LLM, the training of an MLLM undergoes a single-stage or a two-stage process. In both cases, a standard cross-entropy loss is utilized for predicting the next token, serving as an auto-regressive objective.
Single-Stage Training. This possibility is explored by LLaMA-Adapter Gao et al. (2023) which introduces additional trainable parameters to encapsulate the visual knowledge and manage text-only instruction learning at the same time. To achieve this, the model undergoes joint training using image-text pairs and instructions, operating on separate parameters. Concurrently, the model proposed in Koh et al. (2023b) adapts the final loss function by incorporating two contrastive losses for image-text retrieval. During the training, only three linear layers are updated. On a different line, Kosmos-1 Huang et al. (2023b) considers a frozen visual backbone and trains the language model of 1.3B parameters from scratch.
Flamingo Alayrac et al. (2022) and its open source variants Awadalla et al. (2023); Laurençon et al. (2024), instead, train the cross-attention layers and the Perceiver-based component to connect the visual features with the frozen LLM blocks. Additionally, Otter Li et al. (2023b) extends Flamingo’s training to increment its in-context capabilities. Given the amount of training data currently available, approaches like SPHINX-X Gao et al. (2024) opt to perform a single all-in-one training stage in which to update all model components, possibly also using text-only data to preserve the conversational capabilities of the LLM.
Two-Stage Training. In the first of the two training stages, the objective is to align the image features with the text embedding space. After this stage, the outputs tend to be fragmented and not coherent. Therefore, a second step is done to improve multimodal conversational capabilities. LLaVA Liu et al. (2023e, d) is among the first to introduce a visual instruction-following training scheme, which is performed as a second training stage updating the parameters of both the multimodal adapter and LLM. During the first stage, instead, only the multimodal adapter is trainable. Differently, MiniGPT-4 Zhu et al. (2023a) is notable for training solely the linear layer responsible for multimodal alignment across both stages. In the second stage, it uses filtered data, collected and refined through the model itself after the first stage.
Another approach, as demonstrated in InstructBLIP Dai et al. (2023), involves the freezing of the visual encoder and LLM. In both training stages, only the Q-Former and the connection module are trainable. In contrast to previous approaches where the visual backbone remains frozen, mPLUG-Owl Ye et al. (2023c, d) updates it in the initial stage, facilitating the capture of both low- and high-level visual information. Also, in the second stage text-only and multimodal data are used jointly to increase alignment. Differently, Shikra Chen et al. (2023f) updates all weights in both stages, with the only exception of the visual backbone which is kept frozen.
Training Data. During the first (or single) training stage, the datasets predominantly consist of large-scale, publicly available, and uncurated data. For instance, the Conceptual Captions 3M (CC3M) dataset Sharma et al. (2018) is composed of 3M images paired with textual captions specifically designed for image captioning systems. Unlike the widely-used and curated MS-COCO Lin et al. (2014) dataset, which serves similar purposes, images and captions in CC3M are gathered from the web, showcasing a broader spectrum of styles and content. Similarly, the LAION family Schuhmann et al. (2021, 2022) represents an extended collection of non-curated image-text pairs sourced from web pages, providing a rich resource for pre-training multimodal language models. Additionally, the COYO-700M Byeon et al. (2022) dataset stands out as a significant resource, containing 747M image-text pairs. Notably, each alt-text in COYO-700M is linked to an image within HTML documents. Furthermore, DataComp Gadre et al. (2023) presents an extensive pool of 12.8B filtered image-text pairs sourced from common crawl.
It is important to highlight the distinction between datasets used in the initial phase of training, which typically comprise large-scale, uncurated data, and those selected for refinement in subsequent stages. While the former emphasizes diversity and scale, the latter focuses on specificity and task relevance, facilitating a more tailored approach to model optimization. Especially in single-training stage approaches, certain methods Alayrac et al. (2022); Laurençon et al. (2024) also leverage interleaved datasets, which contain images interleaved with text coming from the web, aiming to augment the dataset size for large models Hoffmann et al. (2022). Images within these datasets can be positioned at the beginning or in the middle of a sentence, allowing models to support arbitrarily interleaved sequences of images and text as input, thereby enhancing flexibility in input formats by blending textual and visual elements. Among these datasets, the most used are WebLI Chen et al. (2023j), composed of 10B images and image-text pairs, and MMC4 Zhu et al. (2023d), an extension of the text-only C4 Raffel et al. (2020) dataset composed of 365M documents and 156B tokens relatives to different concepts, and OBELICS Laurençon et al. (2024), an open and curated collection of interleaved image-text web documents, containing 141M documents, 115B text tokens, and 353M images.
In the context of visual instruction tuning, which constitutes the second training stage for MLLMs, the available amount of data is limited. This limitation is mainly due to the creation process which is time-consuming and less well-defined. In this phase, different datasets are used to improve performances on a series of downstream tasks. Among them, LLaVA-Instruct Liu et al. (2023e) is a collection of GPT-4 generated multimodal instruction-following data. It comprises 158k unique language-image descriptions, spanning various types of tasks including 58k conversations, 23k detailed descriptions, and 77k complex reasoning. Similarly, LRV-Instruction Liu et al. (2023c) initially consisted of 400k visual instructions generated by GPT-4, and more recently, it has been updated with an additional set of 300k visual instructions. To enhance robustness in instruction tuning, LRV-Instruct also includes negative instructions organized across three semantic levels, showing that instruct-tuned MLLMs on this dataset suffer less from hallucination compared to the original versions. Moreover, LLaVAR Zhang et al. (2023i) considers publicly available OCR tools to collect results on 422k text-rich images from the LAION dataset. The pipeline first collects 422k noisy text-rich images and then extracts the text through OCR models. With the help of GPT-4, the results and captions are used to create 16k conversations, also including specific questions to create complex instructions which can be helpful to boost performance on new tasks.
Model LLM Visual Encoder Supporting Model Main Tasks & Capabilities
ContextDET Zang et al. (2023) OPT-6.7B Swin-B - Visual Dialogue, VQA, Captioning, Detection, REC, RES
DetGPT Pi et al. (2023) Vicuna-13B EVA ViT-g G-DINO Visual Dialogue, Detection
VisionLLM Wang et al. (2023e) Alpaca-7B Intern-H Deformable-DETR VQA, Captioning, Detection, Segmentation, REC
BuboGPT Zhao et al. (2023c) Vicuna-7B EVA ViT-g RAM, G-DINO, SAM Visual Dialogue, Audio Understanding, Captioning, GroundCap
ChatSpot Zhao et al. (2023b) Vicuna-7B CLIP ViT-L - Visual Dialogue, VQA, Captioning, Referring
GPT4RoI Zhang et al. (2023g) LLaVA-7B OpenCLIP ViT-H - Visual Dialogue, VQA, Captioning, Referring
ASM Wang et al. (2023d) Husky-7B EVA ViT-g - VQA, Captioning, Referring
LISA Lai et al. (2023) LLaVA-13B CLIP ViT-L SAM Visual Dialogue, Captioning, RES
PVIT Chen et al. (2023a) LLaVA-7B CLIP ViT-L RegionCLIP Visual Dialogue, VQA, Captioning, Referring
GLaMM Rasheed et al. (2023) Vicuna-7B OpenCLIP ViT-H SAM Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
Griffon Zhan et al. (2023) LLaVA-13B CLIP ViT-L - REC, Detection, Phrase Grounding
LLaFS Zhu et al. (2023c) CodeLLaMA-7B CLIP RN50 - Few-Shot Segmentation
NExT-Chat Zhang et al. (2023a) Vicuna-7B CLIP ViT-L SAM Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
GSVA Xia et al. (2023b) LLaVA-13B CLIP ViT-L SAM VQA, Segmentation, REC, RES
Lenna Wei et al. (2023) LLaVA-7B CLIP ViT-L G-DINO VQA, Captioning, REC
LISA++ Yang et al. (2023b) LLaVA-13B CLIP ViT-L SAM Visual Dialogue, Captioning, RES
LLaVA-G Zhang et al. (2023d) Vicuna-13B CLIP ViT-L OpenSeeD, S-SAM Visual Dialogue, REC, RES, Grounding
PixelLLM Xu et al. (2023a) FlanT5-XL-3B EVA ViT-L SAM Referring, REC, RES, GroundCap
PixelLM Ren et al. (2023b) LLaVA-7B CLIP ViT-L - Visual Dialogue, RES
VistaLLM Pramanick et al. (2023) Vicuna-13B EVA - Visual Dialogue, VQA, Referring, REC, RES, GroundCap
ChatterBox Tian et al. (2024b) LLaVA-13B CLIP ViT-L iTPN-B, DINO Visual Dialogue, Referring, REC, GroundCap
GELLA Qi et al. (2024) LLaVA-13B CLIP ViT-L Mask2Former Segmentation, RES, GroundCap
PaLI-3 Chen et al. (2023i) UL2-3B SigLIP ViT-g VQ-VAE VQA, Captioning, Retrieval, RES
Table 2: Summary of MLLMs with components specifically designed for visual grounding and region-level understanding. For each model, we indicate the LLM used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM, and any supporting models used to perform the task ($\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with PEFT techniques; $\bigstar$: frozen). Gray color indicates models not publicly available.
3 Tackling Visual Tasks with MLLMs
Standard MLLMs can tackle visual understanding tasks, such as VQA, captioning and multi-turn conversation. However, recently there has been an interest in addressing more fine-grained visual tasks, such as visual grounding and image generation.
3.1 Visual Grounding
The visual grounding capabilities of an MLLM correspond to the ability to carry a dialogue with the user that includes the positioning of the content, also referred to as a referential dialogue Chen et al. (2023f). In particular, You et al. (2023) introduce referring as the ability to understand the content of an input region and can be evaluated on tasks such as region captioning and referring expression generation. Conversely, grounding is associated with localizing regions of a given textual description and corresponds to tasks such as referring expression comprehension (REC), referring expression segmentation (RES), phrase grounding, and grounded captioning. Two main components are required to equip MLLMs with these capabilities: a region-to-sequence method to process input regions and a sequence-to-region method to ground nouns and phrases. A summary of the MLLMs with visual grounding capabilities is reported in Table 2.
Region-as-Text. The most common way to output regions is to directly insert them into generated text as a series of coordinates, represented as numbers or as special tokens dedicated to location bins. Shikra Chen et al. (2023f), Kosmos-2 Peng et al. (2023), MiniGPT-v2 Chen et al. (2023e), Ferret You et al. (2023), CogVLM Wang et al. (2023c), SPHINX Lin et al. (2023b), Qwen-VL Bai et al. (2023b), and Griffon Zhan et al. (2023) convert bounding boxes into text by indicating two points. VisionLLM Wang et al. (2023e), VistaLLM Pramanick et al. (2023), LLaFS Zhu et al. (2023c), and ChatSpot Zhao et al. (2023b) allow the MLLM to handle polygons by representing them as a series of points.
Embedding-as-Region. Another solution is to read input regions through region encoders and provide the output regions as embeddings extracted from the last layer of the MLLM to a decoder. For input regions, GLaMM Rasheed et al. (2023), GPT4RoI Zhang et al. (2023g), ASM Wang et al. (2023d) and ChatterBox Tian et al. (2024b) leverage features of the image encoder to perform ROI align on the bounding box, whereas PVIT Chen et al. (2023a) exploits RegionCLIP Zhong et al. (2022). PixelLLM Xu et al. (2023a) and LLaVA-G Zhang et al. (2023d) use the prompt encoder of SAM Kirillov et al. (2023) and Semantic-SAM Li et al. (2023e) respectively. For output regions, LISA Lai et al. (2023), GLaMM, GSVA Xia et al. (2023b), NeXt-Chat Zhang et al. (2023a), and LISA++ Yang et al. (2023b) send the embedding corresponding to special tokens to the mask decoder of SAM, LLaVA-G to OpenSeeD Zhang et al. (2023c), Lenna Wei et al. (2023) to Grounding-DINO Liu et al. (2023i), and PixelLM Ren et al. (2023b) to a custom lightweight pixel decoder.
Differently, ContextDET Zang et al. (2023) introduces a decoder that receives the latent embedding of the noun with learnable queries, performs a cross-attention with image features, and then uses a segmentation head. ChatterBox Tian et al. (2024b) combines features from the iTPN-B encoder Tian et al. (2023) and the MLLM and provides them to the DINO detector Zhang et al. (2022a). GELLA Qi et al. (2024) presents a fusion module in Mask2Former Cheng et al. (2022) to propose masks based on multi-modal image features and an association module to assign latent embeddings to them. PaLI-3 Chen et al. (2023i) converts embeddings into segmentation masks through a VQ-VAE Van Den Oord et al. (2017) decoder.
Text-to-Grounding. Other approaches are based on open-vocabulary models that accept textual categories as input. DetGPT Pi et al. (2023) generates a list of categories for Grounding-DINO. BuboGPT Zhao et al. (2023c) leverages a combination of RAM, Grounding-DINO, and SAM and matches tags with nouns in the output sequence.
Model LLM Visual Encoder Supporting Model Main Tasks & Capabilities
GILL Koh et al. (2023a) OPT-6.7B CLIP ViT-L SD v1.5 Visual Dialogue, Retrieval, Image Generation
Emu Sun et al. (2023b) LLaMA-13B EVA ViT-g SD v1.5 Visual Dialogue, VQA, Captioning, Image Generation
SEED Ge et al. (2023a) OPT-2.7B EVA ViT-g SD v1.4 VQA, Captioning, Image Generation
DreamLLM Dong et al. (2023) Vicuna-7B CLIP ViT-L SD v2.1 Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
LaVIT Jin et al. (2023) LLaMA-7B EVA ViT-g SD v1.5 VQA, Captioning, Image Generation
MGIE Fu et al. (2024) LLaVA-7B CLIP ViT-L SD v1.5 Image Editing
TextBind Li et al. (2023f) LLaMA-2-7B EVA ViT-g SD XL Visual Dialogue, VQA, Captioning, Image Generation
Kosmos-G Pan et al. (2023) Magneto-1.3B CLIP ViT-L SD v1.5 Image Generation, Compositional Image Generation
MiniGPT-5 Zheng et al. (2023) Vicuna-7B EVA ViT-g SD v2.1 Visual Dialogue, Image Generation, Interleaved Generation
SEED-LLaMA Ge et al. (2023b) LLaMA-2-13B EVA ViT-g SD unCLIP Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
CoDi-2 Tang et al. (2023) LLaMA-2-7B ImageBind SD unCLIP Visual Dialogue, Audio Understanding, Image Generation, Image Editing
Emu2 Sun et al. (2023a) LLaMA-33B EVA ViT-E SD XL Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
LLMGA Xia et al. (2023a) LLaVA-13B CLIP ViT-L SD XL Visual Dialogue, VQA, Image Generation, Image Editing
SmartEdit Huang et al. (2023c) LLaVA-13B CLIP ViT-L SD Image Editing
VL-GPT Zhu et al. (2023b) LLaMA-7B CLIP ViT-L SD v1.5 Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
MM-Interleaved Tian et al. (2024a) Vicuna-13B CLIP ViT-L SD v2.1 VQA, Captioning, REC, Image Generation, Interleaved Generation
JAM Aiello et al. (2024) LLaMA-7B - CM3Leon Image Generation, Interleaved Generation
Table 3: Summary of MLLMs with components specifically designed for image generation and editing. For each model, we indicate the LLM (✻: LLM variants) used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM, and any supporting models used to perform the task ($\lozenge$: training from scratch; $\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with PEFT techniques; $\bigstar$: frozen). Gray color indicates models not publicly available.
3.2 Image Generation and Editing
While initial MLLMs excelled in extracting information from visual data, recent research included the generation of visual outputs. This advancement is realized through integrating MLLMs with image generation mechanisms, predominantly embodied by the Stable Diffusion (SD) Rombach et al. (2022) models. These models feature a denoising U-Net Ronneberger et al. (2015) architecture conditioned on textual or visual embeddings, through cross-attention layers. A complete list of the analyzed models is presented in Table 3.
Connecting MLLMs with Diffusion Models. GILL Koh et al. (2023a) is the pioneer in mapping the output embedding space of a frozen LLM to that of a frozen diffusion model. Specifically, inspired by Q-Former, a mapper component is trained by minimizing the $\ell_{2}$ distance between the image output representation of the language model and the expected conditioning embedding of SD.
While GILL refrains from fine-tuning both the LLM and the diffusion U-Net, alternative methodologies fine-tune the language model to expand its multimodal generation capabilities. In this vein, Kosmos-G Pan et al. (2023) is developed through a training regime that integrates the output of the LLM with an encoder-decoder structure, leveraging a reconstruction loss and the minimization of the distance within a CLIP-text embedding. Similarly, MiniGPT-5 Zheng et al. (2023) includes the reconstruction loss of diffusion models in addition to the alignment loss of GILL. Moreover, it divides the overall training process into two distinct phases: the initial phase concentrates on text-to-image generation, while the subsequent is focused on interleaved vision-and-language generation. Distinctly, researchers have studied the alignment of discrete Jin et al. (2023); Ge et al. (2023a, b) and continuous visual tokens Zhu et al. (2023b) extracted from input images with the SD conditioning embedding. This is usually achieved by fine-tuning the textual model Zhu et al. (2023b); Ge et al. (2023a, b). Conversely, Jin et al. (2023) fine-tune both the LLM and the SD U-Net.
A different approach has been studied by Li et al. (2023f) which proposes to fine-tune the LLM by adding two special tokens (i.e., <start> and <end>), and directly encode the generated text between these two tokens using the text encoder in the SD model. Similarly, in Xia et al. (2023a) the LLM is trained to output detailed language-based generation prompts which are employed for generation or editing tasks. The U-Net is fine-tuned with longer and more detailed textual captions. Furthermore, in DreamLLM Dong et al. (2023) an alignment loss is eschewed in favor of a score distillation loss while keeping the U-Net frozen. Additional research endeavors have been conducted to introduce MLLMs in the field of image editing Fu et al. (2024); Huang et al. (2023c); Tang et al. (2023).
End-to-End Pipelines. A different direction is the development of end-to-end training strategies. Specifically, in Sun et al. (2023b, a) the SD U-Net is directly fine-tuned with the continuous visual embeddings generated by the LLM. Tian et al. (2024a) employ a feature synchronizer, that intervenes in intermediate layers of the LLM and diffusion decoder to cross-attend multi-scale high-resolution image features. Furthermore, end-to-end training approaches have been employed for non-diffusion-based generators, such as VQ-GAN Esser et al. (2021), as demonstrated in the study by Lu et al. (2023a). Differently, Aiello et al. (2024) propose a methodology to mix an LLM architecture with an autoregressive generator, CM3Leon Yu et al. (2023a), via bi-directional cross-attention across the architectures of both models.
3.3 Other Modalities and Applications
Video Understanding. Although much of the research focuses on images, some works propose MLLMs specifically designed to handle video sequences. These models process video frames independently, using CLIP-based backbones to extract frame-level features which are then combined using pooling mechanisms Li et al. (2023j); Maaz et al. (2023) or Q-Former based solutions Li et al. (2023h); Ren et al. (2023a). The connection between visual features and the language model mainly follows the same trend as image-based MLLMs, with linear projections being the most common choice. However, there are also some attempts to develop video-specific adapters Liu et al. (2023g); Ma et al. (2023a) that can capture fine-grained temporal information. In addition to encoding video frames, some works Munasinghe et al. (2023); Zhang et al. (2023b) also employ audio features to enrich the representation of input video sequences. Furthermore, effective strategies for visual instruction tuning are also designed in the video domain Song et al. (2024), enabling more effective understanding of long video sequences.
Any-Modality Models. Almost all models described so far treat a single modality as input to the LLM. However, a significant body of work focuses on designing effective solutions that can handle multiple modalities. This is usually achieved by aligning multimodal features through Transformer blocks such as Q-Former Chen et al. (2023c); Panagopoulou et al. (2023) and Perceiver Zhao et al. (2023d), or by utilizing ImageBind to effectively extract features that are inherently multimodal Su et al. (2023). Images, videos, and audio are the most commonly treated modalities. Additionally, some works also effectively encode 3D data Yin et al. (2023d) and IMU sensor signals Moon et al. (2023). While all these solutions can manage multimodal inputs, approaches like NExT-GPT Wu et al. (2023c) and Unified-IO 2 Lu et al. (2023a) are also capable of generating outputs of different modalities.
Domain-Specific MLLMs. In addition to dealing with generic visual inputs, some research efforts are dedicated to developing MLLMs for specific domains and applications, either training the model starting from a pre-trained LLM or fine-tuning an existing MLLM with domain-specific data. Some examples are MLLMs designed for document analysis and text-intensive visual inputs Lv et al. (2023); Ye et al. (2023a), those proposed for embodied AI and robotics Driess et al. (2023); Mu et al. (2023), and those tailored for specific domains such as medicine Li et al. (2023d) and autonomous driving Xu et al. (2023c). A complete list of domain-specific MLLMs is reported in the supplementary.
4 Conclusion and Future Directions
In this survey, we have provided a comprehensive overview of the recent evolution of MLLMs, first focusing on how to equip LLMs with multimodal capabilities and then exploring the main tasks addressed by these models. Based on the analysis presented, in the following, we outline important open challenges and promising future research directions to further empower MLLMs.
Multimodal Retrieval-Augmented Generation. While retrieval-augmented generation (RAG) is a consolidated technique in LLMs Lewis et al. (2020); Asai et al. (2023), its application in MLLMs is still under-explored. We believe that the emergence of VQA datasets that require external retrieved knowledge Chen et al. (2023k); Mensink et al. (2023) may enable the development of MLLMs with RAG capabilities Hu et al. (2023b); Caffagni et al. (2024).
Correction of Hallucinations. Several studies Liu et al. (2023b); Zhu et al. (2023a) show that MLLMs tend to exhibit high hallucination rates, especially when generating longer captions. While some solutions are emerging to mitigate this problem Liu et al. (2023b); Wang et al. (2023a); Wu et al. (2023d); Yin et al. (2023b); Jing et al. (2023), understanding and correcting the underlying causes of hallucinations remains an important open challenge that is worth addressing to allow the application of these models in more critical contexts (e.g., medicine) and guarantee their accuracy and trustworthiness.
Prevent Harmful and Biased Generation. Ensuring the safety and fairness of large-scale models is of fundamental interest in the community. Recent works show that models trained on web-crawled data are prone to generate inappropriate and biased content. Although recent efforts are being made to reduce this phenomenon in text-to-image generative models Schramowski et al. (2023); Friedrich et al. (2023); Poppi et al. (2024), further exploration is needed to prevent the same behavior in MLLMs Pi et al. (2024).
Reduce Computational Load. As shown in the supplementary, MLLMs are highly computationally demanding. Effective strategies Chu et al. (2024) are needed to reduce computational requirements and enable more accessible development of MLLMs. Possible directions entail reducing training requirements both in terms of model scale and data quantity and optimizing the inference stage.
Limitations
This survey provides a comprehensive review of visual-based MLLMs. Although we have made a significant effort to include all relevant works available to the date of submission, the review might have missed some minor works, and might not have a complete coverage of MLLMs treating modalities that are different from the visual one. Additionally, given the space constraints required by the submission venue, we have restricted our explanations of existing approaches so as to include only the most relevant novelty points. We encourage the reader to refer to the original papers for further technical details and implementation notes.
Acknowledgments
This work has been partially supported by the projects: PNRR-M4C2 (PE00000013) “FAIR - Future Artificial Intelligence Research” funded by the European Commission, the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR) funded by the European Union - NextGenerationEU (CUP B53C22001770006), and the PRIN project “CREATIVE: CRoss-modal understanding and gEnerATIon of Visual and tExtual content” co-funded by the Italian Ministry of University and Research (CUP B87G22000460001).
Appendix A Handling Images of Different Resolutions and Aspect Ratios
Most existing MLLMs perceive images in a low resolution and a fixed squared aspect ratio. Some works Liu et al. (2023d); You et al. (2023); Chen et al. (2023j) have demonstrated that adopting visual backbones trained on higher resolutions leads to fewer hallucinations and improved multimodal understanding abilities, translating into better performance over tasks that require fine-grained details. However, scaling an MLLM to arbitrary input resolutions and aspect ratios raises two important concerns: (i) the adaptation issue of switching from small images seen during training to larger ones at inference time and (ii) computational costs provided by the increased number of tokens in both the visual encoder and the LLM, given by the quadratic complexity of the attention-based architectures. In the following, we distinguish three different approaches to address these problems.
Positional-Encoding Interpolation. These models interpolate the positional encoding of their visual backbones, trained at low resolutions, to handle high-resolution images. While being simple, these methods are prone to adaptation issues. As a consequence, they partially mitigate this issue by performing at least one high-resolution training stage. To reduce the input sequence length to the LLM, and thus the computational cost, MiniGPT-v2 Chen et al. (2023e) and VILA Lin et al. (2023a) propose to project multiple visual tokens together into the same token within the embedding space of the LLM. For the same reason, mPLUG-Owl2 Ye et al. (2023d) and Qwen-VL Bai et al. (2023b) compress the visual features into fixed-length sequences, independent of the resolution, using learnable queries. The latter further saves computation in most layers of the ViT backbone due to a window attention mechanism.
Sub-Images Slicing. To avoid the adaptation issue, some methods propose to slice a high-resolution image into multiple sub-images of fixed size according to the native resolution of their visual encoder. Then, each sub-image is processed independently by the visual backbone, along with the whole image downsized at the same resolution, and the features are concatenated to obtain the global representation. SPHINX Lin et al. (2023b) divides the input image in a squared grid of sub-images (i.e., $2\times 2$ or $3\times 3$) at the training resolution of the visual backbone. Moreover, to handle rectangular aspect ratios, SPHINX pads the image to reach the desired square size. For extreme aspect ratios, the padding leads to sub-images which are only composed of padding. Hence, in SPHINX-X Gao et al. (2024) a skip token is introduced to replace noisy tokens associated with only padding sub-images and reduce the sequence length provided to the LLM, increasing efficiency. Similarly, LLaVA-NeXT Liu et al. (2024b) ignores sub-images composed only of padding and handles different shapes of grids, by introducing a special token that indicates when a row of sub-images ends. Monkey Li et al. (2023l) uses a Perceiver-like resampler to extract fixed-length sequences from each sub-image and trains on an image-text dataset curated by several vision expert models integrated by ChatGPT. InfiMM-HD Liu et al. (2024a) proposes a dynamic resolution adaptation training stage to increase the image size up to 1,344 pixels. It employs gated cross-attention layers (as in Flamingo) to inject the visual features into the LLM, without increasing the input sequence length. LLaVA-UHD Xu et al. (2024) finds the optimal partitioning scheme leading to sub-images that most resemble the native resolution and aspect ratio of the visual encoder. The number of visual tokens is compressed through a Perceiver-like adapter and employs a spatial schema in such a way that the LLM can understand the grid of sub-images.
Others. Another solution, namely OtterHD Li et al. (2023a), can seamlessly deal with any resolution or aspect ratio, as it directly feeds large image patches of $30\times 30$ pixels to the LLM, without the need for a visual encoder. LLaVA-HR Luo et al. (2024), instead, introduces a mixture-of-resolution adaptation to fuse into the ViT layers high-resolution features extracted with a CNN and low-resolution ones produced by the ViT itself.
Appendix B Additional Training Data
Specific training datasets are required to empower MLLMs with visual grounding and image generation capabilities. Here we briefly describe the common choices in this domain.
Visual Grounding. To enable visual grounding, MLLMs can be trained directly on task-specific data using predetermined instruction templates. For instance, CoinIt Pramanick et al. (2023) is a unified set of 14 benchmarks converted into an instruction-tuning format, spanning from single-image coarse-level to multi-image region-level tasks. An additional training step is usually performed on an instruction-tuning dataset, such as LLaVa-Instruct Liu et al. (2023h), to preserve the conversational capabilities of the MLLM. However, some methods create their custom datasets to simultaneously improve the grounding and conversational capabilities. Specifically, Shikra Chen et al. (2023f), DetGPT Pi et al. (2023), ChatSpot Zhao et al. (2023b), and PVIT Chen et al. (2023a) leverage LLMs Achiam et al. (2023); OpenAI (2022) to combine regions and captions from datasets that present both annotations (e.g., COCO). Differently, Kosmos-2 Peng et al. (2023) and Ferret You et al. (2023) exploit an open-vocabulary detector Li et al. (2022) to ground noun chunks parsed from captions and then reconstruct referring expressions. ASM Wang et al. (2023d), GLaMM Rasheed et al. (2023), and LLaVA-G Zhang et al. (2023d) propose automated pipelines comprising multiple steps based on off-the-shelf models for generating large corpora of conversations grounded in their corresponding images.
Image Generation and Editing. To perform image generation, datasets containing both textual captions and images are required, as the one mentioned in Sec. 2.4 (e.g., LAION-400M, COYO-700M, and COCO). To enable interleaved text-image generation, MMC4, OBELICS, and VIST Huang et al. (2016) are popular choices. Instead, for image editing tasks, additional datasets like the one introduced in InstructPix2Pix Brooks et al. (2023) and MagicBrush Zhang et al. (2023e) are typically used.
VQA Captioning MLLM Evaluation
[…]
Flamingo Alayrac et al. (2022) 82.0 - 65.7 - 57.1 138.1 75.4 - - - - - - -
BLIP-2 Li et al. (2023g) 65.0 41.0 19.6 61.0 42.5 144.5 - 85.3 1293.8 - 46.4 38.1 22.4 -
OpenFlamingo Awadalla et al. (2023) 52.7 - 27.5 - 24.2 75.9 59.5 - - - - - - -
MiniGPT-4 Zhu et al. (2023a) 53.7 32.2 - - - - - - 581.7 23.0 42.8 45.1 22.1 23.1
mPLUG-Owl Ye et al. (2023c) 59.5 40.9 - - - - - - 967.3 46.6 34.0 - - -
ChatBridge Zhao et al. (2023d) - 41.8 - - - - 82.5 - - - - - - -
InstructBLIP Dai et al. (2023) 69.4 49.5 33.4 63.1 50.7 102.2 82.8 78.9 1212.8 36.0 53.4 58.2 25.6 25.3
Shikra Chen et al. (2023f) 77.4 - - - - 117.5 - - - 58.8 - 79.9 - -
Emu Sun et al. (2023b) 62.0 46.0 38.3 - - 117.7 - - - - - - 36.3 -
SVIT Zhao et al. (2023a) 80.3 64.1 56.4 70.0 60.8 - - - 1565.8 69.1 61.9 - - -
BLIVA Hu et al. (2024) - - 42.9 - 58.0 - 87.1 - 1669.2 - - - - -
IDEFICS Laurençon et al. (2024) 60.0 45.2 36.0 - 30.9 - - - - 54.5 - - -
Qwen-VL Bai et al. (2023b) 78.2 57.5 38.9 68.2 61.5 120.2 81.0 - 1487.6 60.6 58.2 56.7 - -
DreamLLM Dong et al. (2023) 56.6 - 38.1 - 34.9 115.4 - - - 49.9 - - 35.9 -
LLaVA-1.5 Liu et al. (2023d) 80.0 63.3 53.6 71.6 61.3 - - 85.9 1531.3 67.7 61.6 70.7 35.4 23.6
CogVLM Wang et al. (2023c) 82.3 - - - - 148.7 94.9 87.9 - 77.6 72.5 77.8 51.1 34.5
LION Chen et al. (2023d) - 51.6 - - - 139.3 87.1 88.9 - - - - - -
mPLUG-Owl2 Ye et al. (2023d) 79.4 56.1 54.5 - - 137.3 - 86.2 1450.2 64.5 57.8 25.0 36.2 25.3
SPHINX Lin et al. (2023b) 80.2 62.9 46.8 69.1 - - - 90.8 1560.2 67.1 71.6 74.3 36.6 27.5
Emu2 Sun et al. (2023a) 84.9 65.1 54.9 - 66.6 - - - - - 62.8 - 48.5 -
Honeybee Cha et al. (2023) - - - - - - - - 1632.0 73.6 68.6 77.5 - -
Unified-IO 2 Lu et al. (2023a) 79.4 - - 88.7 - 125.4 - 87.7 - 71.5 61.8 - - -
VILA Lin et al. (2023a) 80.8 63.3 60.6 73.7 66.6 115.7 74.2 84.2 1570.1 70.3 62.8 73.0 38.8
SPHINX-X Gao et al. (2024) 81.1 63.8 61.9 74.5 - - - 89.6 1485.3 71.3 73.0 70.2 40.9 42.7
Table 4: Performance analysis on 14 evaluation benchmarks for VQA, image captioning, and MLLM evaluation. Best scores are in bold, second best are underlined.
Appendix C Evaluation
MLLMs are evaluated across different benchmarks, taking into account both more classic visual comprehension and recognition skills and advanced multimodal conversation capabilities. Table 4 shows the performance of the most common MLLMs on both standard VQA and captioning datasets and benchmarks specifically designed for evaluating MLLMs. In the following, we detail the datasets reported in the table and other benchmarks typically used for the evaluation of MLLMs.
C.1 Standard Benchmarks
One of the most important skills of MLLMs is their ability to effectively answer questions based on the given input image. This ability is quantitatively evaluated across several visual question-answering datasets, measuring the accuracy (Antol et al., 2015) of the answers provided by the MLLM.
VQAv2 Goyal et al. (2017) is an extended and balanced version of VQA (Antol et al., 2015) built by collecting similar images for the same question, but whose answer is different compared to the original one. This makes it difficult to perform favorably for those models that ignore visual information and only rely on language priors while answering questions. The reported results are related to the test-dev split.
GQA Hudson and Manning (2019) is based on Visual Genome scene graph annotations (Krishna et al., 2017) and comprises 113k images and 22M questions focusing on scene understanding and compositionality. We report the results over the test split, which contains 10% of the total images.
OKVQA Marino et al. (2019) is a benchmark to study how vision-and-language models can address visual questions whose answers cannot be completely found in the image, encouraging systems that also rely on external knowledge. The test set has 14,055 open-ended questions.
VizWiz Gurari et al. (2018) originates from authentic situations involving individuals with visual impairments who have taken images and articulated accompanying inquiries about them, together with 10 responses. The validation split consists of 4,319 images paired with their corresponding questions, while the test split encompasses roughly 8,000 instances.
ScienceQA (SQA) Lu et al. (2022) evaluates models over challenging multimodal multiple-choice questions about 3 subjects (i.e., natural science, language science, and social science), 26 topics, 127 categories, and 379 skills. Each question is annotated with explanations linked to relevant lectures. The test set includes 4,241 examples.
Visual Spatial Reasoning (VSR) Liu et al. (2023a) contains images from COCO, each paired with a caption mentioning two concepts and the spatial relation between them. Models have to choose if a given caption is true or false according to the picture. MLLMs are typically evaluated on the 616 samples from the zero-shot test split.
IconQA Lu et al. (2021) tests the visual reasoning abilities of vision-and-language models on three types of questions: multiple-image-choice, multiple-text-choice, and fill-in-the-blank. The dataset stems from real-world problems found in math textbooks and focuses on abstract images (i.e., icons). There are 107,439 questions, 20% of which makes up for the test split.
${}^{\text{T}}$) Singh et al. (2019) is a dataset based on pictures from Open Images Kuznetsova et al. (2020) and challenges OCR capabilities of vision-and-language models. The test set comprises 5,734 examples.
OCR-VQA Mishra et al. (2019) presents a new task in visual question answering by interpreting text within images and involves a collection of 207,572 images of book covers, accompanied by more than 1M question-answer pairs.
Comprehensively describing the visual input is another important skill desired in MLLMs. To evaluate this, various image captioning datasets are commonly employed. As regards the evaluation metric, the CIDEr score Vedantam et al. (2015), which is the reference metric for the task, is used to compare generated image descriptions with ground-truth captions.
COCO Lin et al. (2014) contains more than 120k images, each associated with five human-generated captions. For captioning tasks, the splits defined by Karpathy and Fei-Fei (2015) are typically employed, with 113k, 5k, and 5k images respectively for train, validation and test.
Flickr30k Young et al. (2014) comprises 31,783 images, depicting diverse everyday activities, events, and scenes. Complementing these images are 158,915 captions, obtained through crowd-sourcing techniques.
nocaps Agrawal et al. (2019) represents a benchmark for novel object captioning, boasting an extensive collection of almost 400 novel object categories compared to the COCO dataset. The validation and test sets include approximately 4.5k and 10.6k images, obtained from Open Images Kuznetsova et al. (2020). Each image is annotated with 11 human-generated captions. Both validation and test sets are further categorized into in-domain, near-domain, and out-of-domain, where images from the out-of-domain subset contain object categories that are never present in COCO.
TextCaps Sidorov et al. (2020) includes 145k captions aligned with 28k images. The goal is to recognize and understand the text in images and provide an effective caption that describes the entire visual content. This requires the model to possess OCR capabilities along with image description skills.
C.2 MLLM-Specific Benchmarks
Thoroughly evaluating MLLMs is challenging and remains an open frontier. While evaluating on standard datasets represents a valid choice, many benchmarks designed for MLLMs have been recently proposed. They require very strong perception and cognitive skills to succeed, and often they query for deep domain-specific knowledge. To facilitate the evaluation, many works propose to leverage state-of-the-art proprietary models (e.g., ChatGPT (OpenAI, 2022), GPT-4 (Achiam et al., 2023)) to automatically judge candidate answers. In Table 4, we report the performance of some models on a subset of these new benchmarks.
POPE Li et al. (2023k) is a valuable benchmark for evaluating object hallucination challenges within MLLMs. This dataset encompasses several distinct subsets, namely random, popular, and adversarial, which are generated utilizing a variety of sampling methodologies. Cumulatively, it is a binary classification query dataset that comprises 8,910 entries, facilitating comprehensive investigations into the phenomenon of object hallucination within the context of MLLMs.
MME Fu et al. (2023) is an evaluation benchmark that aims to assess proficiency in various communication modalities through 14 tasks covering comprehension and manipulation across modalities like quantification, spatial determination, color identification, and others.
MMBench (MMB) Liu et al. (2023k) includes approximately 3,000 multiple-choice questions, distributed across 20 distinct domains. Questions are curated to evaluate the efficacy of MLLM across diverse task paradigms. These competencies are systematically arranged into a hierarchical taxonomy, delineating overarching categories such as perception and reasoning, while also outlining granular capabilities including object localization and attribute inference.
SEED-Bench (SEED) Li et al. (2023c) is specifically designed to evaluate LLMs and MLLMs across 12 dimensions spanning from scene understanding to OCR and action recognition. The benchmark consists of 19k multiple-choice questions written by human annotators.
[…]
MM-Vet Yu et al. (2023b) evaluates MLLMs over 16 tasks covering six fundamental vision-and-language capabilities such as recognition, OCR, knowledge, language generation, spatial awareness, and math. The benchmark comprises 200 images and 218 questions. The evaluation scores are obtained from GPT-4 by few-shot prompting.
[…]
MMMU Yue et al. (2023) is a challenging benchmark targeting domain-specific knowledge of multimodal models. It consists of 10.5k test samples drawn from university textbooks or online courses spanning six main disciplines. Questions may contain multiple images interleaved with text. Exact matching and word matching are used to assess the correctness of an answer for multiple-choice and open-ended questions respectively. Models are evaluated on zero or few-shot settings.
Tiny LVLM Shao et al. (2023) focuses on six multimodal capabilities distributed among 2.1k image-question pairs. It introduces a new evaluation metric called ChatGPT ensemble evaluation (CEE). In practice, given the question and the ground-truth solution, ChatGPT is queried with five different prompts to assign the candidate answer either 0 or 1, and the scores are eventually ensembled.
TouchStone Bai et al. (2023c) is a visual dialog benchmark with manually annotated open-world images, totaling 908 questions corresponding to five major categories of abilities and 27 sub-tasks. The evaluation score is computed by an LLM such as GPT-4, which is asked to compare a candidate answer with a reference one. The latter is computed by GPT-4 itself, with fine-grained annotations of the query image being part of the prompt.
C.3 Visual Grounding Evaluation
The assessment of visual grounding capabilities of MLLMs comprises a variety of standard referring tasks, including region captioning, referring expression generation (REG), and region-level question answering, as well as grounding tasks like referring expression comprehension (REC), referring expression segmentation (RES) and grounded captioning. As regards evaluation metrics, for REC the accuracy is computed by assuming as correct predictions the ones that correspond to an intersection over union with the ground-truth above 0.5 (Acc@0.5). For referring expression segmentation the cumulative intersection over union (cIoU) is considered, while for region captioning METEOR Banerjee and Lavie (2005) and CIDEr Vedantam et al. (2015) are commonly used. However, few methods introduce their own benchmarks to evaluate the performance in more realistic scenarios, with grounded conversations that may involve multiple rounds. Quantitative results on the REC, RES, and region captioning tasks are respectively reported in Table 5, Table 6, and Table 7.
RefCOCO RefCOCO+ RefCOCOg
Model val testA testB val testA testB val(U) test(U)
Kosmos-2 Peng et al. (2023) 52.3 57.4 47.3 45.5 50.7 42.2 60.6 61.7
Shikra Chen et al. (2023f) 87.8 91.1 81.8 82.9 87.8 74.4 82.6 83.2
Qwen-VL Bai et al. (2023b) 88.6 92.3 84.5 82.8 88.6 76.8 86.0 86.3
Ferret You et al. (2023) 89.5 92.4 84.4 82.8 88.1 75.2 85.8 86.3
MiniGPT-v2 Chen et al. (2023e) 88.7 91.7 85.3 80.0 85.1 74.5 84.4 84.7
CogVLM Wang et al. (2023c) 92.8 94.8 89.0 88.7 92.9 83.4 89.8 90.8
Griffon Zhan et al. (2023) 90.1 93.4 86.1 84.8 90.5 77.8 86.1 87.2
LION Chen et al. (2023d) 89.8 93.0 85.6 84.0 89.2 78.1 85.5 85.7
NExT-Chat Zhang et al. (2023a) 85.5 90.0 77.9 77.2 84.5 68.0 80.1 79.8
SPHINX Lin et al. (2023b) 91.0 92.7 86.6 86.6 91.1 80.4 88.2 88.4
Lenna Wei et al. (2023) 90.3 93.2 87.0 88.1 90.1 84.0 90.3 90.3
LLaVA-G Zhang et al. (2023d) 89.2 - - 81.7 - - 84.8 -
Unified-IO 2 Lu et al. (2023a) 90.7 - - 83.1 - - 86.6 -
MM-Interleaved Tian et al. (2024a) 89.9 92.6 86.5 83.0 88.6 77.1 85.2 84.9
SPHINX-X Gao et al. (2024) 90.6 93.7 86.9 85.5 90.5 79.9 88.3 88.5
Table 5: Performance analysis on the RefCOCO benchmarks for referring expression comprehension (REC). Best scores are in bold, second best are underlined.
RefCOCO and RefCOCO+ Mao et al. (2016) are collections of referring expressions based on images from the COCO dataset. They were gathered through the ReferItGame Kazemzadeh et al. (2014), a two-player game where the first player examines an image featuring a segmented target object and formulates a natural language description referring to that object. The second player, who has access only to the image and the referring expression, selects the corresponding object. Players swap roles if they perform correctly, otherwise they receive a new object and image for description. The RefCOCO dataset has no constraints on the natural language and consists of 142,209 expressions for 50,000 objects across 19,994 images. Instead, in the RefCOCO+ players are disallowed from using location words in their referring expressions and it has 141,564 expressions for 49,856 objects in 19,992 images. Evaluation is performed on 1,500, 750, and 750 images corresponding to the validation, testA, and testB splits for both datasets.
RefCOCOg Yu et al. (2016) was collected by a set of annotators who wrote natural language referring expressions for objects in COCO images, and another set of annotators who selected objects corresponding to given referring expressions. When a selected object was correct, the corresponding referring expression was inserted in the dataset. It consists of 85,474 referring expressions for 54,822 objects in 26,711 images. Evaluation is carried out on 1,300 and 2,600 images corresponding to the validation and test splits.
Visual Genome Krishna et al. (2017) connects structured image concepts to language and comprises 108,077 images along with detailed descriptions of all objects present in them, providing 5.4M region descriptions and 1.7M visual question-answer pairs. This dataset is typically used for region-level captioning and question-answering.
Visual7W Zhu et al. (2016) is a visual question-answering dataset that combines textual descriptions with image regions through object-level grounding. It comprises 328k question-answer pairs on 47k COCO images, together with 1.3M human-generated multiple-choice and more than 560k object groundings from 36,579 categories.
RefCOCO RefCOCO+ RefCOCOg
Model val testA testB val testA testB val(U) test(U)
LISA Lai et al. (2023) 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
GLaMM Rasheed et al. (2023) 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9
NExT-Chat Zhang et al. (2023a) 74.7 78.9 69.5 65.1 71.9 56.7 67.0 67.0
GSVA Xia et al. (2023b) 79.2 81.7 77.1 70.3 73.8 63.6 75.7 77.0
LLaVA-G Zhang et al. (2023d) 77.1 - - 68.8 - - 71.5 -
PixelLLM Xu et al. (2023a) 76.9 78.5 74.4 69.2 72.1 64.5 70.7 72.4
GELLA Qi et al. (2024) 76.7 80.5 73.6 67.0 73.2 60.6 70.4 71.5
Table 6: Performance analysis on the RefCOCO benchmarks for referring expression segmentation (RES). Best scores are in bold, second best are underlined.
RefCOCO Visual Genome
Model METEOR CIDEr METEOR CIDEr
Kosmos-2 Peng et al. (2023) 14.1 62.3 - -
GPT4RoI Zhang et al. (2023g) - - 17.4 145.2
ASM Wang et al. (2023d) 20.8 103.0 18.0 145.1
GLaMM Rasheed et al. (2023) 16.2 106.0 19.7 180.5
NExT-Chat Zhang et al. (2023a) 13.6 79.6 - -
PixelLLM Xu et al. (2023a) 14.3 82.3 19.9 148.9
Table 7: Performance analysis on the RefCOCO and Visual Genome benchmarks for region captioning. Best scores are in bold, second best are underlined.
GRIT Peng et al. (2023) is a large-scale dataset of grounded image-text pairs (i.e., noun phrases or referring expressions associated with regions of the image) based on a subset of COYO-700M and LAION-2B. The construction pipeline consists of two steps: (i) extracting noun chunks from the captions and grounding them to bounding boxes with an open-vocabulary detector (e.g., GLIP); (ii) expanding the noun chunks to referring expressions by exploiting their dependency relations in the original caption. The resulting dataset comprises 91M images, 115M text spans, and 137M associated bounding boxes.
ReasonSeg Lai et al. (2023) is a benchmark introduced for the reasoning segmentation task, which consists of providing segmentation masks for complex and implicit query texts. Images are from OpenImages Kuznetsova et al. (2020) and ScanNetv2 Dai et al. (2017) and are annotated with text instructions and corresponding segmentation masks. The resulting dataset comprises 1,218 image-instruction pairs. Evaluation metrics are the same as the RES standard benchmark. Two extended variants, ReasonDet Wei et al. (2023) and ReasonSeg-Inst Yang et al. (2023b), are respectively introduced for reasoning detection and reasoning instance segmentation tasks.
Grounding-anything Dataset (GranD) Rasheed et al. (2023) is a dataset designed for the grounded conversation generation (GCG) task, which aims to construct image-level captions with phrases associated with segmentation masks in the image. This dataset was built with an automated annotation pipeline composed of four stages: (i) object localization with the corresponding semantic label, segmentation mask, attributes, and depth information, (ii) extracting relationships between detected objects, (iii) combining previously collected relations to produce dense captions, (iv) enriching captions with contextual information. It comprises annotations for 11M SAM Kirillov et al. (2023) images. Another dataset, $\text{GranD}_{f}$, is introduced for further fine-tuning and evaluating over the GCG task. It was gathered by extending Flickr30k Young et al. (2014), RefCOCOg, and PSG Yang et al. (2022) through GPT-4 and by manually annotating a set of samples. It comprises 214k image-grounded text pairs with 2.5k validation and 5k test samples. Evaluation metrics include METEOR and CIDEr for captioning, class-agnostic mask AP for grounding, intersection over union for segmentation, and mask recall for grounded captioning.
Grounded-Bench Zhang et al. (2023d) is a benchmark introduced to assess the capabilities of an MLLM in carrying a grounded visual chat. It is built on top of the LLaVA-Bench Liu et al. (2023h), comprising conversational data generated with GPT-4 and instance annotations from COCO. It is expanded using 1,000 images with 7,000 entities from COCO annotated through an automated pipeline that involves GPT-4 to associate noun phrases from captions to ground-truth instances.
MUSE Ren et al. (2023b) is a multi-target reasoning segmentation dataset. It was created with an automated pipeline on top of 910k instance segmentation masks from the LVIS dataset Gupta et al. (2019) by exploiting GPT-4V to combine instance categories with natural language descriptions. The resulting dataset comprises 246k question-answer pairs, averaging 3.7 targets per answer.
COCO
Model FID CLIP-I CLIP-T
Stable Diffusion Rombach et al. (2022) 9.22 0.667 0.302
Stable Diffusion XL Podell et al. (2023) - 0.674 0.310
GILL Koh et al. (2023a) 12.20 0.684 -
Emu Sun et al. (2023b) 11.66 0.656 0.286
SEED Ge et al. (2023a) - 0.682 -
DreamLLM Dong et al. (2023) 8.46 - -
LaVIT Jin et al. (2023) 7.40 - -
NExT-GPT Wu et al. (2023c) 11.28 - -
Kosmos-G Pan et al. (2023) 10.99 - -
SEED-LLaMa Ge et al. (2023b) - 0.707 -
Emu2 Sun et al. (2023a) - 0.686 0.297
VL-GPT Zhu et al. (2023b) 11.53 - -
Unified-IO 2 Lu et al. (2023a) 13.39 - -
MM-Interleaved Tian et al. (2024a) 7.90 - -
Table 8: Image generation results on the COCO dataset. Best scores are in bold, second best are underlined.
MagicBrush
Model DINO CLIP-I CLIP-T
InstructPix2Pix Brooks et al. (2023) 0.698 0.854 0.292
MagicBrush Zhang et al. (2023e) 0.868 0.934 0.302
MGIE Fu et al. (2024) 0.903 0.943 0.317
SmartEdit Huang et al. (2023c) 0.815 0.914 0.305
Table 9: Image editing results on the MagicBrush benchmark.
ChatterBox-300k Tian et al. (2024b) is a benchmark established to evaluate models on multimodal dialogue systems in multi-round referring and grounding. The dataset is built on images from Visual Genome Krishna et al. (2017) providing bounding boxes, object relationships, and object attributes information to GPT-4 to generate question-answer pairs.
C.4 Image Generation and Editing Evaluation
To evaluate image generation and editing results, a set of different benchmarks is usually utilized. In terms of evaluation metrics, Fréchet Inception Distance (FID) Heusel et al. (2017) is the reference metric to evaluate generated images. It quantitatively assesses the congruence between the distribution of synthetically generated images and the distribution of real ones. A diminution in the FID score indicates an enhanced alignment between the two distributions, denoting a superior visual quality and realism within the generated images.
Other metrics measure the coherence of the generated image with the input prompt and the real ground-truth image corresponding to it. Specifically, CLIP-I and DINO scores consist of computing the cosine similarity between generated and ground-truth images leveraging CLIP Radford et al. (2021) and DINO Caron et al. (2021) as visual backbones. CLIP-T, instead, measures image-text alignment through cosine similarity between input captions and generated images, using CLIP to encode both images and textual prompts.
DreamBench
Model DINO CLIP-I CLIP-T
DreamBooth Ruiz et al. (2023) 0.668 0.803 0.305
Kosmos-G Pan et al. (2023) 0.694 0.847 0.287
CoDi-2 Tang et al. (2023) 0.703 0.852 0.311
Emu2 Sun et al. (2023a) 0.766 0.850 0.287
Table 10: Subject-driven image generation results on the DreamBench dataset.
COCO is employed for evaluating text-to-image generation. The evaluation is conducted using either the original validation set comprising 41k samples or a subset of 30k samples randomly selected from the same set. Results on this dataset of MLLMs with image generation capabilities are reported in Table 8.
VIST Huang et al. (2016) is specifically curated for the task of interleaved image-text generation. It includes 34k and 5k samples for training and evaluation. Each sample is a sequence consisting of 5 images accompanied by 5 textual narratives that collectively form a coherent story.
MagicBrush Zhang et al. (2023e) is a benchmark in the area of image editing and contains a collection of 10,000 manually annotated triplets, each consisting of a source image, an editing instruction, and the corresponding target image. Performances on this benchmark are reported in Table 9.
DreamBench Ruiz et al. (2023) is a benchmark that evaluates the generative capabilities of the models on subject-driven generation. Specifically, it contains 30 subjects, each illustrated with 4 to 6 images, and 25 template prompts enabling modification and accessorization of the given subjects. Results on this benchmark are shown in Table 10.
Model Hardware Type #
FlamingoAlayrac et al. (2022) TPUv4 1,535
PaLI Chen et al. (2023j) TPUv4 1,024
IDEFICS Laurençon et al. (2024) A100 512
SPHINX Lin et al. (2023b) A100 32
Emu Sun et al. (2023b) A100 128
VILA Lin et al. (2023a) A100 128
BLIP-2 Li et al. (2023g) A100 16
SEED-LLaMA Ge et al. (2023b) A100 64
Shikra Chen et al. (2023f) A100 8
MiniGPT-v2 Chen et al. (2023e) A100 8
InstructBLIP Dai et al. (2023) A100 16
BLIVA Hu et al. (2024) A6000 8
CleverFlamingo Chen et al. (2023b) A100 8
LLaVA 1.5 Liu et al. (2023d) A100 8
LLaVA Liu et al. (2023e) A100 8
MiniGPT-4 Zhu et al. (2023a) A100 4
FROMAGe Koh et al. (2023b) A100 1
LaVIN Luo et al. (2023) A100 8
Table 11: Summary of the hardware required to train common MLLMs.
Appendix D Computational Requirements
To provide a quantification of the computational requirements necessary to train an MLLM, we compare some of the most common models in Table 11 and indicate for each of them the type and number of GPUs/TPUs employed during training. Except for Flamingo and PaLI, which are trained on a large amount of TPUs, all other models employ A100 or A6000 GPUs. As it can be seen, most MLLMs distribute training across 8 A100s.
Moreover, in Figure 2 we show for each MLLM the total amount of GPU training hours, approximating 1 TPU hour as 1.5 GPU hours. Notably, models like Flamingo, PaLI, and IDEFICS require a significant amount of GPU time (in the order of magnitude of a few hundred thousand GPU hours). Instead, lighter models like LLaVA only require a few hundred GPU hours to complete training.
Figure 2: Number of GPU training hours for various MLLMs. Here 1 TPU hour is approximated as 1.5 GPU hours following public benchmarks.
Model LLM Visual Encoder Main Tasks & Capabilities
VideoChat Li et al. (2023h) StableVicuna-13B EVA ViT-g Visual Dialogue, VQA, Captioning
Video-ChatGPT Maaz et al. (2023) Vicuna-7B CLIP ViT-L Visual Dialogue, VQA, Captioning
Video-LLaMA Zhang et al. (2023b) Vicuna-7B EVA ViT-g Visual Dialogue, Captioning, VQA, Audio Understanding
BT-Adapter Liu et al. (2023g) Vicuna-7B CLIP ViT-L Visual Dialogue, Captioning, VQA, Retrieval
LLaMA-VID Li et al. (2023j) Vicuna-13B EVA ViT-g Visual Dialogue, VQA, Captioning
PG-Video-LLaVA Munasinghe et al. (2023) LLaVA-1.5-13B CLIP ViT-L Visual Dialogue, Captioning, VQA, Grounding
TimeChat Ren et al. (2023a) LLaMA-2-7B EVA ViT-g Visual Dialogue, Captioning, Temporal Grounding, Highlight Detection
Vista-LLaMA Ma et al. (2023a) Vicuna-7B EVA ViT-g Visual Dialogue, VQA, Captioning
Table 12: Summary of video-based MLLMs. For each model, we indicate the LLM used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM ($\bigstar$: frozen LLM; $\blacklozenge$: LLM fine-tuning; $\blacktriangle$: LLM fine-tuning with PEFT techniques).
Appendix E Additional Details on Other Modalities and Applications
Video Understanding. As a complement of Sec. 3.3, we report in Table 12 a summary of the main characteristics of video-based MLLMs. For each model, we indicate the LLM used as starting point, which in some cases is initialized with the parameters of a pre-trained MLLM, the visual encoder, and the main tasks and capabilities of the MLLM. Additionally, we specify whether the LLM is kept frozen, is entirely fine-tuned, or is fine-tuned with PEFT-based strategies.
3D Understanding. MLLMs are also applied to 3D data for solving complex tasks like 3D VQA, 3D conversation, and 3D dense captioning. Differently from standard visual encodings which exploit 2D pre-trained embeddings, in the context of 3D data, appropriate strategies are designed to project them to the LLM space. In 3D-LLM Hong et al. (2023), 3D scenes are rendered in different views and 3D features are built using an EVA-CLIP backbone connected to a fine-tuned BLIP-2 model. Similarly, Xu et al. (2023b) employ a pre-trained Point-BERT Yu et al. (2022) as 3D encoder and conducts a two-stage training that initially aligns the input features via an MLP projection layer, and then performs an instruction tuning phase of the model. Differently, in Point-Bind Guo et al. (2023), 3D point-clouds are aligned with ImageBind Girdhar et al. (2023) and by leveraging I2P-MAE Zhang et al. (2023f) as 3D encoder. This alignment allows the introduction of new tasks such as any-to-3D generation and 3D embedding-space arithmetic. Recently, LL3DA Chen et al. (2023g) introduces the Interactor3D module, which consists of a frozen 3D scene encoder, a visual prompt encoder, and a Q-Former to address 3D captioning and VQA.
Model LLM Visual Encoder Main Tasks & Capabilities
Document Analysis
mPLUG-DocOwl Ye et al. (2023a) mPLUG-Owl-7B CLIP ViT-L Visual Dialogue, Captioning, VQA
Kosmos-2.5 Lv et al. (2023) Magneto-1.3B Pix2Struct ViT-L Text Recognition, Image-to-Markdown Generation
UReader Ye et al. (2023b) mPLUG-Owl-7B CLIP VIT-L Visual Dialogue, VQA, Captioning, Information Extraction
mPLUG-PaperOwl Hu et al. (2023a) mPLUG-Owl-7B CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
LLaMA-SciTune Horawalavithana et al. (2023) LLaMA-13B CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
DocPedia Feng et al. (2023) Vicuna-7B Swin-B Visual Dialogue, VQA, Information Extraction
Embodied AI
EmbodiedGPT Mu et al. (2023) LLaMA-7B EVA ViT/g, RN50 Visual Dialogue, VQA, Captioning, Task Planning
PaLM-E Driess et al. (2023) PaLM-540B ViT-22B Visual Dialogue, VQA, Captioning, Task Planning, Manipulation
Medical Vision Learning
PMC-VQA Zhang et al. (2023h) PMC-LLaMA-7B PMC-CLIP RN50 VQA
LLaVA-Med Li et al. (2023d) LLaVA-7B CLIP ViT-L Visual Dialogue, VQA
Qilin-Med-VL Liu et al. (2023f) CN-LLaMA2-13B CLIP ViT-L Visual Dialogue, VQA
Autonomous Driving
Dolphins Ma et al. (2023b) OpenFlamingo-7B CLIP ViT-L Visual Dialogue, VQA, Captioning, Traffic Condition Understanding
DriveGPT4 Xu et al. (2023c) LLaMA-2-7B CLIP ViT-L Visual Dialogue, VQA, Captioning
Food Understanding
FoodLLM Yin et al. (2023c) LISA-7B CLIP ViT-L Visual Dialogue, VQA, Nutrition Estimation, RES
Table 13: Summary of MLLMs designed for domain-specific applications. For each model, we indicate the LLM used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM ($\bigstar$: frozen LLM; $\blacklozenge$: LLM fine-tuning; $\blacktriangle$: LLM fine-tuning with PEFT techniques). Gray color indicates models not publicly available.
Any-Modality Models. Several studies focus on extending the reasoning capabilities of the MLLMs by including multiple modalities, such as video, 3D, and audio. A line of research investigates the usage of dedicated pathways to input the different modalities to the LLM. UniVAL Shukor et al. (2023) maps the features from each modality encoder into the shared representation space of the LLM through dedicated linear projections. X-LLM Chen et al. (2023c) leverages Q-Former interfaces for the image and video modalities, interpreting the video as a sequence of independent frames, each one encoded as an image. For the speech modality, it uses a C-Former interface that compresses the feature sequence from the speech encoder into token-level embeddings. X-InstructBLIP Panagopoulou et al. (2023) and ChatBridge Zhao et al. (2023d) propose to freeze both the modality encoders and the LLM and to leverage, respectively, dedicated Q-Former or Perceiver adapters for each modality. To maximize feature compatibility, AnyMAL Moon et al. (2023) uses an encoder that has already been aligned to a text embedding space for each modality, including also IMU signals, and a dedicated adapter, which is a Perceiver for the visual modality and linear layers for the others. On the other hand, PandaGPT Su et al. (2023) and NExT-GPT Wu et al. (2023c) exploit a single frozen multimodal encoder (i.e., ImageBind) to extract features from different modalities. OneLLM Han et al. (2024) builds a unified universal encoder and a universal projection module by mixing multiple image projection modules and a modality router to align input signals with language. CAT Ye et al. (2024) adds a clue aggregator to aggregate question-aware audio-visual hidden features and produce clue tokens that are provided to the LLM.
In addition to handling different modalities in input to the LLM, some works investigate the generation of outputs of different modalities. For example, NExT-GPT Wu et al. (2023c) introduces signal tokens in the LLM that indicate whether the diffusion-based decoder for a specific modality has to be activated. Moreover, the signal tokens are provided to a transformer-based output projector to condition the generation. Similarly, M2UGen Hussain et al. (2023) handles the music modality by using the LLM output corresponding to signal tokens, along with unimodal music features from a music encoder, to condition the generation of an audio encoder. LLMBind Zhu et al. (2024) indicates the conditioning text to generate image, video, or audio by wrapping it in special tokens. Thus, this text is provided to the corresponding modality-specific diffusion model. Unified-IO2 Lu et al. (2023a) uses VQ-GAN decoders for both image and audio modalities to decode output discrete tokens and can generate surface normals, depth, and segmentation masks for the input images. AnyGPT Zhan et al. (2024) interprets all the continuous non-text modalities as discrete tokens in both input and output, using, respectively, multimodal tokenizers and de-tokenizers. To enable the 3D modality, LAMM Yin et al. (2023d) introduces a novel instruction tuning dataset and benchmark that comprise both image-text and point cloud-text instruction-response pairs, covering a wide range of 2D and 3D tasks.
Interactive and Compositional Systems. A different trend is to build systems that can combine multiple tools (i.e., existing vision-only or vision-and-language models), usually through ChatGPT or another LLM. In particular, these approaches aim to let the user interact with the LLM which is in charge of selecting the useful tools to carry out complex tasks. In this context, some solutions study how to prompt ChatGPT Wu et al. (2023a); Yang et al. (2023c) to invoke visual foundation models. GPT4Tools Yang et al. (2023a), instead, employs open-source LLMs such as LLaMA and OPT, that are fine-tuned with PEFT techniques to use tools for performing a wide range of visual tasks. Differently, Liu et al. (2023l) introduce more sophisticated user-chatbot interactions, through the incorporation of mouse-based pointing instructions on images or videos.
While in all these approaches the LLM does not directly handle the visual input which is instead processed by other external tools, in LLaVA-Plus Liu et al. (2023h) the query image is directly input to the MLLM (i.e., LLaVA) and is therefore involved during the selection and invocation of the most helpful tool according to the user needs. This is achieved also thanks to the introduction of a new instruction-following use tool dataset, which is employed to fine-tune the MLLM.
Domain-Specific MLLMs. Finally, in Table 13 we summarize the main characteristics of domain-specific MLLMs, also in this case indicating for each model the LLM used as starting point.